OpenStudy (anonymous):

how about this now, another stor problem margie is certain that every time she parks in the municipal it cost her at least $2.50. if the garage charges 40 cents plus 30 cents for each half hour how long is morgie's car parked?

OpenStudy (anonymous):

\[.40+.60x=2.50\]

OpenStudy (anonymous):

no that is not right.

OpenStudy (anonymous):

it is a penny a minute, so if you want time in minutes convert to pennies and get \[40+x=250\]

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

and work with pennies rather than dollars so we can ignore the stupid decimal

OpenStudy (anonymous):

30 cents for half an hour. half an hour is 30 minutes, so parking is 1 cent per minute

OpenStudy (anonymous):

every minute cost a penny, and you also have to pay 40 cents on top of that. so if x is the number of minutes you park, you have to pay \[x+40\] cents. for example if you park 30 minutes (half an hour) you pay 30+40=70 cents

OpenStudy (anonymous):

now you know that you pay at least 250 cents. so put \[x+40=250\] and solve for x

OpenStudy (anonymous):

in one step. \[x+40=250\] \[x=250-40=210\]

OpenStudy (anonymous):

and so if you park for 210 minutes you pay 250 cents or $2.50

OpenStudy (anonymous):

210 minutes is three and a half hours. so if you park for 3 and a half hours you pay $2.50